OpenStudy (anonymous):

Are these identities correct (trigonometry) I would be learning them , so i want to verify

OpenStudy (anonymous):

sin A + sin B = 2 sin ½ (A + B) cos ½ (A − B) sin A − sin B = 2 sin ½ (A − B) cos ½ (A + B) cos A + cos B = 2 cos ½ (A + B) cos ½ (A − B) cos A − cos B = −2 sin ½ (A + B) sin ½ (A − B)

hartnn (hartnn):

or http://www.purplemath.com/modules/idents.htm